# Moduli spaces of untwisted wild Riemann surfaces

### v1

@@ 4.2.4. The fundamental group
4.2.7 Remark.
  1. (1)
    If $\mathsf{M}$ is the analytic stack attached to a manifold $X$, then the transformations mentioned in the previous remark are always trivial, hence Definition 4.2.5 recovers the usual fundamental group of $(X,x)$.
  2. (2)
    There is an alternative way to define $\pi_{1}(\mathsf{M}^{top},x)$ only involving fundamental groups of topological spaces: choose an atlas $X\rightarrow\mathsf{M}$ and consider the topological groupoid $X_{\bullet}^{top}=R^{top}\rightrightarrows X^{top}$. Consider the classifying space $BX_{\bullet}$ of $X_{\bullet}$, as defined in [Noo12, §4.1]. It comes with a natural map $BX_{\bullet}\rightarrow\mathsf{M}$ [Noo12, Proposition 6.1]; one defines $\pi_{1}(\mathsf{M},x)=\pi_{1}(BX_{\bullet},\tilde{x})$ for a point $\tilde{x}$ above $x$. The fact that this definition is independent of choices and coincides with the one given above follows from [Noo12, proofs of Theorem 10.5, Theorem 6.3].
4.2.8. Example
The approach described in the previous remark allows us to describe fundamental groups of quotient stacks, cf. [Noo12, Remark 4.3]. Firstly, if $\mathsf{M}=[*/G]$ for a (discrete) group $G$ then $X=*$ and $BX_{\bullet}=EG/G$, where $EG$ is the universal principal $G$-bundle (in particular, it is a simply connected space on which $G$ acts freely), so $\pi_{1}(\mathsf{M},*)=G$. More generally, if $G$ acts on a topological space $X$ and $\mathsf{M}$ is the quotient stack $[X/G]$ (i.e. it is attached to the groupoid $X_{\bullet}=G\times X\rightrightarrows X$ whose maps are the projection on $X$ and the $G$-action) then $BX_{\bullet}$ is the quotient $X\times_{G}EG$ of $X\times EG$ by the antidiagonal action of $G$ (Borel construction). In particular, if $X$ is simply connected then $\pi_{1}([X/G],x)=G$ (no matter how big the stabilisers of points of $X$ are). For instance, with the notation of section 4.1.9, the fundamental group of $\mathsf{M}_{g,m}$ is $\Gamma_{g,m}$.

### v2

@@ A.2.4. The fundamental group
A.2.8. Example
The approach described in the previous remark allows us to describe fundamental groups of quotient stacks, cf. [Noo12, Remark 4.3]. Firstly, if $\mathsf{M}=[*/G]$ for a (discrete) group $G$ then X=X=* and $BX_{\bullet}=EG/G$, where $EG$ is the universal principal $G$-bundle (in particular, it is a simply connected space on which $G$ acts freely), so $\pi_{1}(\mathsf{M},*)=G$. More generally, if $G$ acts on a topological space $X$ and $\mathsf{M}$ is the quotient stack $[X/G]$ (i.e. it is attached to the groupoid $X_{\bullet}=G\times X\rightrightarrows X$ whose maps are the projection on $X$ and the $G$-action) then $BX_{\bullet}$ is the quotient $X\times_{G}EG$ of $X\times EG$ by the antidiagonal action of $G$ (Borel construction).
In particular, if $X$ is simply connected then $\pi_{1}([X/G],x)=G$ (regardless of the stabilisers of points of $X$). For instance, with the notation of section A.1.9, the fundamental group of $\mathsf{M}_{g,m}$ is $\Gamma_{g,m}$.